OpenStudy (anonymous):

Just finished my math test on got 21/22 but i dont understand why i got the 1 wrong. It says; What makes this a true statement 9^2= ? I answered -1/18 because i had no idea and thought the answer was 0.09. It says the answer is 1/18, why is this? I will medal the best answer! Just please HELP!

OpenStudy (anonymous):

@e.mccormick @love10129151 can you help?

OpenStudy (anonymous):

Are you sure it said: $$9^2=?$$Because \(9^2=81.\)

OpenStudy (anonymous):

it said 9\[^{-2}\] sorry typo

OpenStudy (anonymous):

\[*9^{-2}\]

OpenStudy (anonymous):

Ah, ok. Use this rule: $$a^{-k}=\frac{1}{a^k}.$$

OpenStudy (anonymous):

So $$9^{-2}=\frac{1}{9^2}=\frac{1}{81}$$

OpenStudy (e.mccormick):

yah, as nerdguy2535 pointed out, it looks like you just transposed the numbers in the answer.

OpenStudy (e.mccormick):

Or perhaps that is another typo and the issue is the - ?

OpenStudy (anonymous):

no i thought the answer was actually 0.09 scientific notation. @nerdguy2535 is right, thanks though!

OpenStudy (e.mccormick):

Kk. Long as you understand what nerdguy2535 did you should be able to answer these. It is a basic thing of how negative exponents work.

OpenStudy (triciaal):

square of a number is always positive when you take the square root 2 solutions

OpenStudy (triciaal):

What makes this a true statement 9^2= ? what were the choices?